# A Statistical Learning Approach for Feature-Aware Task-to-Core Allocation in Heterogeneous Platforms

### v1

@@ 5 Related Work
Probabilistic Methods for DVFS and Task-to-Core Allocation.
@@
Estimating probabilistic worst-case execution time (pWCET) and worst-case energy consumption (pWCEC) in embedded real-time systems has been explored via measurement-based and static approaches Cazorla et al. (2019); Davis and Cucu-Grosjean (2019); Reghenzani et al. (2020); Pallister et al. (2017). Pallister et al. Pallister et al. (2017) analyzed the impact of different instructions on pWCEC, while the extreme value theorem (EVT) Edgar and Burns (2001) predicted upper bounds on performance metrics. However, most existing probabilistic models do not systematically evaluate the statistical importance of system and performance metrics for bounding energy or latency. Our work introduces a statistical learning approach that prioritizes feature significance, enabling low-energy designs and enhanced environment modeling under DVFS and task-to-core allocation.
Statistical Learning.
Researchers have applied statistical learning to identify influential features for energy optimization and scheduling, often based on hardware events or application parameters Sasaki et al. (2007); Cazorla et al. (2019); Liu et al. (2021). Sasaki et al. Sasaki et al. (2007) used decision trees to minimize table look-up overhead in DVFS settings, and Cazorla et al. Cazorla et al. (2019) examined hardware counters to reduce energy consumption. Liu et al. Liu et al. (2021) ranked compiler-generated features by correlation with latency. However, these studies overlooked the potential of runtime performance metrics, sampling techniques, and accuracy trade-offs. We address these gaps by comprehensively evaluating feature correlations and devising a feature selection strategy suitable for parallel scheduling and environment modeling.
Low-Energy DVFS and Task-to-Core Allocation.
A broad body of work investigates low-energy scheduling on multicore platforms using DVFS and task-to-core assignment Xie et al. (2021, 2017); Zhu et al. (2004); Jiang et al. (2019); Chen et al. (2018); Zhou et al. (2018); Kim and Wu (2020); Dinakarrao et al. (2019); Shen et al. (2012); Wang et al. (2017). Xie et al. Xie et al. (2021) reviewed heuristic, meta-heuristic, and machine learning algorithms for parallel scheduling under energy constraints. Despite these advances, many machine learning methods for low-energy DVFS management demand large datasets and incur high computational overhead. Our statistical learning approach alleviates these drawbacks by reducing data requirements while preserving prediction accuracy, making it practical for embedded and heterogeneous processing applications.
Few-Shot RL.
@@
Few-shot learning techniques—including transfer learning, meta-learning, and data augmentation—target reduced data collection in processor scheduling Wang et al. (2020); Lee et al. (2020); Wang et al. (2016); Florensa et al. (2017); Arora and Doshi (2021). Model-agnostic meta-learning (MAML) Finn et al. (2017) adapts to new tasks with minimal data, and model-based reinforcement learning (RL) Moerland et al. (2023) approximates transition functions to reduce reliance on real-world samples. Prior studies Lin et al. (2023); Kim et al. (2021); Zhou and Lin (2021); Zhang et al. (2024) have not fully explored statistical resampling or feature importance evaluations for energy-efficient task-to-core scheduling. Our approach integrates these elements, bridging the gap by combining statistical resampling with a robust feature selection methodology.
Feature Evaluation.
Predictive models for thermal-aware scheduling have been proposed in numerous works Yan et al. (2003); Brooks et al. (2007); Maity et al. (2022); Hosseinimotlagh and Kim (2019); Singla et al. (2015); Li and Wu (2012); Kassab et al. (2021), often relying on utilization or temperature monitoring to avoid thermal throttling Maity et al. (2022); Lin et al. (2023); Kim et al. (2021). Although these models forecast thermal behavior based on transient and ambient states Hosseinimotlagh et al. (2021), they lack a rigorous statistical correlation analysis for feature selection. We address this limitation with a systematic approach that examines feature interdependencies and their effects on energy efficiency and performance, thereby optimizing DVFS strategies with more precise task-to-core allocation.

### v2

@@ 5 Related Work
Probabilistic Methods for DVFS and Task-to-Core Allocation
@@
Probabilistic worst-case execution time (pWCET) and worst-case energy consumption (pWCEC) estimation in embedded real-time systems leverage measurement-based and static methods Cazorla et al. (2019); Davis and Cucu-Grosjean (2019); Reghenzani et al. (2020); Pallister et al. (2017). Pallister et al. Pallister et al. (2017) studied instruction-specific impacts on pWCEC, and extreme value theorem (EVT) Edgar and Burns (2001) provided upper bounds for performance metrics. However, these approaches rarely quantify the statistical significance of system metrics for energy or latency bounds. Our work introduces a hybrid statistical learning framework, systematically prioritizing feature importance to optimize energy-efficient DVFS and task-to-core allocation across diverse platforms, surpassing traditional probabilistic bounds.
Statistical Learning
Statistical learning has been employed to pinpoint features for energy optimization and scheduling, often using hardware events or application traits Sasaki et al. (2007); Cazorla et al. (2019); Liu et al. (2021). Sasaki et al. Sasaki et al. (2007) applied decision trees to streamline DVFS table lookups, while Cazorla et al. Cazorla et al. (2019) used hardware counters for energy reduction. Liu et al. Liu et al. (2021) correlated compiler features with latency. Yet, these efforts underexplored runtime metrics and sampling strategies critical for accuracy. We enhance this by integrating runtime performance metrics with a novel feature selection pipeline, validated across Intel Xeon and Core i7 platforms, improving parallel scheduling and environment modeling precision.
Low-Energy DVFS and Task-to-Core Allocation
Low-energy multicore scheduling via DVFS and task-to-core mapping is well-studied Xie et al. (2021, 2017); Zhu et al. (2004); Jiang et al. (2019); Chen et al. (2018); Zhou et al. (2018); Kim and Wu (2020); Dinakarrao et al. (2019); Shen et al. (2012); Wang et al. (2017). Xie et al. Xie et al. (2021) surveyed heuristic and machine learning methods for energy-constrained scheduling. Many such ML approaches, however, require extensive datasets and computational resources, limiting practicality. Our method mitigates these issues with a lightweight statistical model, reducing data needs while maintaining accuracy, as demonstrated on embedded and heterogeneous systems. Hierarchical multi-agent approaches Pivezhandi et al. (2026b) and graph-driven performance models Pivezhandi et al. (2026a) have also shown promise for DAG workloads, offering scalable alternatives to existing high-overhead techniques.
Few-Shot RL
@@
Few-shot learning, including transfer learning and meta-learning, minimizes data demands in scheduling Wang et al. (2020); Lee et al. (2020); Wang et al. (2016); Florensa et al. (2017); Arora and Doshi (2021). MAML Finn et al. (2017) adapts to new tasks with few samples, and model-based RL Moerland et al. (2023) approximates dynamics to limit real-world data use. Works like Lin et al. (2023); Kim et al. (2021); Zhou and Lin (2021); Zhang et al. (2024) underutilize statistical resampling and feature ranking for energy-efficient scheduling. Recent advances include flow-augmented data generation for few-shot RL Pivezhandi and Saifullah (2026a) and zero-shot LLM-guided resource allocation Pivezhandi and Saifullah (2026b). Our approach fills this gap, combining resampling with a robust feature selection strategy, validated across platforms, to enhance few-shot task-to-core allocation efficiency.
Feature Evaluation
Thermal-aware scheduling models often predict behavior using utilization or temperature data to prevent throttling Yan et al. (2003); Brooks et al. (2007); Maity et al. (2022); Hosseinimotlagh and Kim (2019); Singla et al. (2015); Li and Wu (2012); Kassab et al. (2021). Studies like Maity et al. (2022); Lin et al. (2023); Kim et al. (2021) model transients and ambient conditions Hosseinimotlagh et al. (2021), but lack rigorous statistical analysis of feature correlations. We address this by applying a systematic correlation-based feature evaluation, rigorously tested on real hardware, to uncover interdependencies and optimize DVFS and allocation, improving energy and performance outcomes over prior heuristic-driven methods.
@@ 6 Conclusion
ACKNOWLEDGEMENT
The work was supported by the US National Science Foundation through grants CNS-2301757, CAREER- 2306486, CNS-2306745, and by the US Office of Naval Research through grant N00014-23-1-2151.